# Leveraging Vision-Language Models for Visual Grounding and Analysis of Automotive UI

### v1

@@ Abstract
Modern automotive infotainment systems require intelligent and adaptive solutions to handle frequent User Interface (UI) updates and diverse design variations. We introduce a vision-language framework for understanding and interacting with automotive infotainment systems, enabling seamless adaptation across different UI designs. To further support research in this field, we release AutomotiveUI-Bench-4K, an open-source dataset of 998 images with 4,208 annotations. Additionally, we present a synthetic data pipeline to generate training data. We fine-tune a Molmo-7B-based model using Low-Rank Adaptation (LoRa) and incorporating reasoning generated by our pipeline, along with visual grounding and evaluation capabilities. The fine-tuned Evaluative Large Action Model (ELAM) achieves strong performance on AutomotiveUI-Bench-4K (model and dataset are available on Hugging Face) and demonstrating strong cross-domain generalization, including a +5.2% improvement on ScreenSpot over the baseline model. Notably, our approach achieves 80.4% average accuracy on ScreenSpot, closely matching or even surpassing specialized models for desktop, mobile, and web, such as ShowUI, despite being trained for the infotainment domain. This research investigates how data collection and subsequent fine-tuning can lead to AI-driven progress within automotive UI understanding and interaction. The applied method is cost-efficient and fine-tuned models can be deployed on consumer-grade GPUs.
Key Words: Graphical User Interface, Large Action Model, Vision-Language Model, Automotive UI, Fine-Tuning.
@@ 1 Introduction
Automotive infotainment systems are rapidly evolving, characterized by increasing complexity, dynamic interfaces, and personalization [20]. With manufacturers frequently pushing over-the-air updates and diverse designs proliferating across models, these systems demand intelligent and adaptive solutions capable of handling significant variations [36]. This constant evolution necessitates intelligent systems that can dynamically interpret the visual layout and semantic meaning of interfaces, moving beyond reliance on fixed structural assumptions. Visual Language Models offer a promising approach by integrating computer vision with natural language understanding, enabling systems to interpret visual information and user intent in a more human-like manner [17]. While these models have demonstrated success in understanding and interacting with interfaces in domains like desktop, mobile and web [3, 10], their application to the distinct environment of automotive infotainment systems remains significantly under-explored. This represents a research gap, given the unique demands and rapid evolution of in-vehicle interfaces. Applying VLMs effectively within the automotive context introduces non-trivial challenges. Infotainment Human-Machine Interfaces exhibit vast heterogeneity across car models and manufacturers, featuring custom icons, menus, and interaction paradigms [20]. Therefore, a robust model must generalize effectively across disparate screen layouts and styles while maintaining high precision in understanding on-screen elements and their context. Moreover, sophisticated interaction often requires reasoning about the HMI state, extending beyond simple object detection or element localization.
To bridge this gap, we introduce in this work a highly adaptable vision-language framework designed for understanding and interacting with automotive infotainment systems, enabling seamless adaptation across different UI designs. Our key contributions include:
@@
    A synthetic data generation pipeline developed to enhance model performance and generalization for small VLMs (7B or less) through parameter efficient fine-tuning.
@@
Through empirical evaluation, we demonstrate the efficacy of our proposed framework. The fine-tuned model establishes a new performance benchmark on our AutomotiveUI-Bench-4K dataset and exhibits remarkable adaptability, significantly improving results on the cross-domain ScreenSpot task (+5.2%). Strikingly, its overall accuracy (80.4% average) competes favorably with, and in some cases surpasses, models specifically designed for narrower domains like desktop, mobile, or web (e.g., ShowUI [18]) as shown in Table 4. The robustness of the model is highlighted by its performance even with a restricted training dataset, both in terms of size and domain. It is important to note, however, that the automotive infotainment domain itself encompasses a diverse range of systems and functionalities. The following sections delve into the system architecture, the curation of our novel dataset, and the comprehensive experimental validation, collectively showcasing the significant potential of AI-driven methods for advancing automotive HMI understanding and interaction.
2 Research Scope
The validation of automotive infotainment HMI traditionally relies on established methodologies such as specification-based testing, which verifies HMI functionality against documented requirements in a black-box manner [14], and Hardware-in-the-Loop (HiL) testing, where the infotainment unit operates in a simulated vehicle environment [15, 16]. Within HiL setups, HMI validation often employs visual techniques like Optical Character Recognition (OCR) and template matching to assess the Graphical User Interface (GUI) based on captured screen images [21, 28]. However, these conventional methods, particularly their dependence on OCR and template matching for visual verification, struggle with the complexities of modern, dynamic UI. They exhibit fragility towards visual variations arising from updates, themes, or different languages [26, 5], and lack a deeper semantic understanding of the interface elements and their context [7]. The resulting high maintenance overhead and limited adaptability motivate the need for more intelligent solutions. Therefore, this research focuses on leveraging VLMs to overcome these limitations, aiming to establish a more robust, flexible, and semantically aware framework for understanding and validating automotive HMI. Notably, this work concentrates on the development and evaluation of the VLM-based framework itself and does not provide specific recommendations or methodologies for its integration into production-ready HMI/HiL testing environments.
@@ 3 Method
This chapter outlines our methodology for adapting our framework to automotive infotainment UIs. We discuss the integration of VLMs for tasks like image captioning and visual question answering, and address challenges in adapting these models for automotive UI. We introduce our approach, including synthetic data generation and fine-tuning of baseline models, to develop a specialized VLM for accurate visual grounding and validation.
3.1 VLM Integration
VLMs have emerged as a pivotal area of research, bridging the gap between visual and textual information processing. These models integrate visual capabilities with Large Language Models to handle complex tasks such as image captioning, visual question answering, and multimodal dialogue systems. Recent advancements in VLMs have led to their classification into three main categories based on their input-processing and output-generation capabilities: Vision-Language Understanding Models, Multimodal Input Text Generation Models, and Multimodal Input-Multimodal Output Models [9].
The evolution of VLMs has also seen the development of models specializing in GUI interaction. Recently the research community established the term Large Action Model (LAM) [29]. Notable comprehensive surveys [37, 22, 31] address critical research questions regarding the development of LAMs tailored for GUI tasks, and the evaluation metrics as well as benchmarks such as ScreenSpot [3], necessary to assess their effectiveness. These studies identify that a specialized VLM that is capable of accurate visual grounding is crucial to perform well in various GUI domains such as desktop, mobile, and web [37].
3.2 Domain Adaptation Challenges
As discussed in the previous section, extensive research covers LAMs and GUI agents, which focus on interactions and implicitly derive evaluative functions by inferring necessary actions. In our targeted application of validating requirements via test cases defined in natural language, it is necessary to explicitly evaluate expected outcomes. Small fine-tuned models (7B or less) such as TinyClick [24] or UGround [10] gathered UI data from mobile, desktop, and web applications to fine-tune baseline models such as Qwen2-VL [30] for visual grounding. This raises the following challenges:
@@ 3.2 Domain Adaptation Challenges
    Adapt VLMs for the domain of automotive UI: Modern infotainment systems contain many different UI designs depending on brand, platform, mode of driving and display hardware properties. Icons pose a particular challenge as there is little overlap to generic UIs from desktop, mobile, or web applications.
@@
    Assert evaluation capability: Prior work does not consider explicit evaluation. In order to evaluate requirements or specifications it is necessary to prompt expected results which are visually grounded and evaluated.
@@
    • Original Equipment Manufacturers often have strict information security requirements which makes leveraging proprietary models such as OpenAI GPT-4o or Google Gemini Pro 2.0 challenging due to concerns about data privacy and potential leakage of sensitive information to external servers.
    • Another challenge arises from fine-tuning such models. OpenAI and Google provide fine-tuning services for training and deployment. Details of the training process and the fine-tuned model itself remain closed source.
    • The discontinuation of closed source models entail challenging validation and prompt adaptation iterations in order to ensure the intended functionality.
    • Running model on local servers or even consumer-grade GPUs is highly desirable and necessary to assure a cost-effective solution.
To address these challenges, we introduce ELAM-7B, that goes beyond existing LAMs by integrating evaluation capabilities. This allows the model to conduct test cases. The model is trained with synthetic data generated by our accompanying pipeline.
@@ 3.3 Synthetic Data Pipeline
Object-level captioning for infotainment screens presents significant challenges due to the necessity of incorporating functional, positional, and visual attributes to uniquely identify each element. As previously mentioned, differentiating between Test Actions and Expected Results is required. Both classes necessitate object-level captions that describe either the command initiating element actuation or a statement that shall be evaluated for the current screen. Furthermore, for Expected Results, a binary status (passed or failed) must be assigned, reflecting the validity of the expectation within the image. Empirically, human annotators exhibit a tendency to under-represent failed expectations, likely due to the cognitive burden of formulating incorrect expectations for a displayed infotainment screenshot. Therefore, a hybrid approach, leveraging human expertise for UI element selection via bounding boxes combined with automated caption generation by large teacher models, (e.g., GPT-4o, Gemini 2.0 Flash Thinking), offers a practical balance between annotation quality and time investment. Automated UI element extraction methods like OmniParser V1.5 [19] and a custom YOLOv8 [13] fine-tuned to automotive UI images were found to be insufficiently reliable, frequently resulting in misaligned bounding boxes that negatively impacted the correctness of generated captions.
The following synthetic data pipeline (Fig. 1) was used to train ELAM-7B. Bounding boxes from the dataset were prompted using the Set-of-Mark (SoM) technique [35]. Inspired from UGround [10], we used a bounding box pointed to by an arrow. For Test Actions, prompts were designed to short yet comprehensive instructions that unambiguously map to a single UI element (Fig. 2(a)). Prior to generating each test action, the model was instructed to provide a reasoning detailing the process of identifying the target UI element, including its semantic, positional, and visual characteristics, as well as any relevant relationships to other elements, such as the current context or list headings (Fig. 2).
@@
[…]
To improve the generation of Expected Results, the process was modified to include a prior successful test action leading to the current menu or screen state. Specifically, the model first generated a prior test action, followed by reasoning for the expectation and its “passed/failed” conclusion. Initially, this approach exhibited the same bias as human annotators and generated mostly passed validations. This issue was mitigated by explicitly prompting for evaluations that entail a failed result (Fig. 2(b)). After manually reviewing a small set of samples for GPT-4o, Gemini 2.0 Flash Thinking and Claude 3.5 Sonnet, we decided to use Gemini (gemini-2.0-flash-thinking-exp-01-21) to generate the data for the training procedure. The detailed prompts used for generating the training datasets are displayed in Sections A.1.1, A.1.2 and A.1.3.
To mitigate potential overfitting on the specific linguistic style of the teacher model, we rephrased the text for reasoning, test actions, and expectations using a smaller model (gpt-4o-mini-2024-07-18). This rephrasing step is crucial for ensuring robustness and generalizability when processing inputs from real-world test case data. The teacher model, configured with a temperature of zero to minimize hallucinations, exhibited a tendency towards repetitive use of verbs and sentence structures, necessitating this diversification strategy. The contribution of rephrasing is explored in Table 5.

### v3

@@ Abstract
Modern automotive infotainment systems necessitate intelligent and adaptive solutions to manage frequent User Interface (UI) updates and diverse design variations. This work introduces a vision-language framework to facilitate the understanding of and interaction with automotive UIs, enabling seamless adaptation across different UI designs. To support research in this field, AutomotiveUI-Bench-4K, an open-source dataset comprising 998 images with 4,208 annotations, is also released. Additionally, a data pipeline for generating training data is presented.
A Molmo-7B-based model is fine-tuned using Low-Rank Adaptation (LoRa), incorporating generated reasoning along with visual grounding and evaluation capabilities. The fine-tuned Evaluative Large Action Model (ELAM) achieves strong performance on AutomotiveUI-Bench-4K (model and dataset are available on Hugging Face). The approach demonstrates strong cross-domain generalization, including a +5.6% improvement on ScreenSpot over the baseline model. An average accuracy of 80.8% is achieved on ScreenSpot, closely matching or surpassing specialized models for desktop, mobile, and web, despite being trained primarily on the automotive domain. This research investigates how data collection and subsequent fine-tuning can lead to AI-driven advancements in automotive UI understanding and interaction. The applied method is cost-efficient, and fine-tuned models can be deployed on consumer-grade GPUs.
@@ 1 Introduction
Automotive infotainment systems are rapidly evolving, characterized by increasing complexity, dynamic interfaces, and personalization [23]. With manufacturers frequently deploying over-the-air updates and diverse User Interface (UI) designs becoming prevalent across models, these systems demand intelligent and adaptive solutions capable of handling significant variations [40]. This constant evolution necessitates intelligent systems capable of dynamically interpreting the visual layout and semantic meaning of interfaces, moving beyond reliance on fixed structural assumptions for UI validation.
Visual Language Models offer a promising approach by integrating computer vision with natural language understanding, enabling systems to interpret visual information and user intent in a more human-like manner [19]. While these models have demonstrated success in understanding and interacting with interfaces in domains such as desktop, mobile, and web [4, 10], their application to the distinct environment of automotive infotainment systems remains significantly underexplored. This represents a research gap, given the unique demands and rapid evolution of in-vehicle interfaces. Applying VLMs effectively within the automotive context introduces non-trivial challenges. Automotive UIs exhibit vast heterogeneity across car models and manufacturers, featuring custom icons, menus, and interaction paradigms [23]. Therefore, a robust model must generalize effectively across disparate screen layouts and styles while maintaining high precision in understanding on-screen elements and their context. Moreover, sophisticated interaction often requires reasoning about the UI state, extending beyond simple object detection or element localization.
To bridge this gap, a highly adaptable vision-language framework is introduced, designed for understanding, interacting, and validating automotive infotainment systems and enabling seamless adaptation across different UI designs. The key contributions of this work include:
@@
    A synthetic data generation pipeline developed to enhance model performance and generalization for small VLMs (7B or less) through parameter-efficient fine-tuning.
@@
Through empirical evaluation, the efficacy of the proposed framework is demonstrated. The fine-tuned model establishes a new performance benchmark on the AutomotiveUI-Bench-4K dataset and exhibits remarkable adaptability, significantly improving results on the cross-domain ScreenSpot task (+5.6%). Its overall accuracy (80.8% average) competes favorably with, and in some cases surpasses, models specifically designed for non-automotive domains (e.g., ShowUI [20]), as shown in Table 3. The robustness of the model is highlighted by its performance even with a restricted training dataset, both in terms of size and domain. It is important to note, however, that the automotive infotainment domain itself encompasses a diverse range of systems and functionalities. The following sections cover the system architecture, the curation of the novel dataset, and the comprehensive experimental validation, collectively showcasing the significant potential of AI-driven methods for advancing automotive UI understanding and interaction.
2 Related Work
This section reviews the literature relevant to automated UI interaction and validation. It begins by examining the current landscape of VLMs and the emerging field of Large Action Models applied to general UI understanding. Subsequently, existing domain-specific datasets and common model adaptation techniques are analyzed, highlighting their limitations for automotive applications. The section concludes by discussing traditional validation methodologies to contextualize and define the specific research gaps addressed in this paper.
2.1 Vision Language Models and UI Understanding
VLMs have emerged as a significant area of research, bridging the gap between visual and textual information processing. These models integrate visual capabilities with Large Language Models to handle complex tasks such as image captioning, visual question answering, and multimodal dialogue systems [9]. Key contributions in this field include models like CLIP [28], which demonstrates strong performance in zero-shot image classification by aligning visual and textual embeddings, and BLIP-2 [17], which introduces an efficient pre-training strategy using frozen image encoders and LLMs. Other notable models include Flamingo [1], which exhibits significant few-shot learning capabilities, and LLaVA [21], which enhances LLMs for multimodal understanding through visual instruction tuning.
A significant advancement in this domain is the Molmo series [5], which introduces approaches to multimodal understanding with notable capabilities in spatial reasoning and accurate localization. Unlike traditional VLMs that primarily rely on discrete vision encoders, Molmo models demonstrate enhanced capability in understanding spatial relationships and providing localization through their pointing mechanism. This pointing capability allows models to not only identify UI elements but also provide exact coordinate information for their locations, representing a significant development for UI interaction tasks. The Molmo series is trained on the comprehensive PixMo dataset [5], which contains millions of image-text pairs with spatial annotations, including pointing data that enables models to ground textual descriptions to specific normalized pixel coordinates in images.
As VLMs evolved, models specializing in UI interaction were developed, leading to the establishment of the term Large Action Model (LAM) [33]. Comprehensive surveys [41, 25, 35] have addressed critical research questions concerning the development of LAMs for UI tasks. It has been highlighted that a specialized VLM capable of accurate visual grounding is crucial for strong performance. Frameworks like ShowUI [20] and SeeClick [4] leverage VLMs to map natural language instructions to UI element interactions. Similarly, LAMs such as OS-Atlas [37] extend VLMs by incorporating action generation capabilities suitable for agentic UI navigation tasks. Furthermore, smaller fine-tuned models like TinyClick [27] (based on Microsoft’s Florence-2-Base [38]) and UGround [10] (based on Qwen2-VL [34]) have been developed specifically for UI tasks, often processing UI images and interpreting actions to generate corresponding point coordinates.
The Molmo series differentiates itself from other UI-focused models through its training methodology that incorporates both general multimodal understanding and specialized pointing capabilities. While most LAMs are trained primarily on UI-specific datasets, Molmo models benefit from the diverse PixMo dataset, which includes a smaller but significant portion of UI-related data from the AndroidControl [18] dataset alongside general visual content. This approach enables better generalization capabilities while maintaining strong performance on UI tasks. The pointing mechanism in Molmo models provides more accurate spatial understanding compared to traditional region-based approaches, such as those relying on bounding boxes, used in other models.
However, these models are typically trained on general-purpose interfaces from desktop, mobile, and web domains. This limits their applicability to automotive UIs, which exhibit domain-specific design patterns and custom iconography. Notably, they lack explicit mechanisms for evaluating UI states (e.g., verifying if a seat belt warning is displayed), an essential requirement for automotive validation. While the Molmo series demonstrates enhanced spatial reasoning capabilities through its pointing mechanism, it still faces the same domain adaptation challenges when applied to automotive interfaces with their unique visual characteristics and functional requirements.
2.2 Domain-Specific UI Datasets and Model Adaptation
Existing UI datasets, including AMEX [2] and Android In the Wild [29], focus on mobile or web environments. These datasets primarily annotate UI elements with interaction labels (e.g., “tap the settings icon”) but rarely include evaluative statements. The PixMo dataset [5], while comprehensive in its multimodal coverage with millions of image-text pairs and innovative pointing annotations, also follows this pattern. Its UI-related subset focuses primarily on interaction rather than validation tasks. The dataset’s strength lies in its rich spatial annotations that enable accurate localization, but it lacks the evaluation-centric labels necessary for automotive UI validation scenarios.
In contrast, the proposed AutomotiveUI-Bench-4K introduces a dual-label structure (Test Action and Expected Result) to simulate real-world validation scenarios, aligning with the need for compliance checks. This approach differs fundamentally from existing datasets, including PixMo. While these datasets excel in interaction modeling, they do not address the critical evaluation dimension required for automotive UI testing.
Parameter-efficient fine-tuning techniques such as Low-Rank Adaptation (LoRa) [12] have been used to adapt VLMs for UI tasks, as seen in UGround [10]. However, these approaches focus solely on interaction and neglect evaluation. Likewise, prior synthetic data pipelines [10, 27] generate interaction-focused annotations. These adaptation methods are extended in this work by the introduction of a novel pipeline. This pipeline balances Test Actions with Expected Results to train evaluation-aware models.
2.3 Traditional Automotive UI Validation and Research Gaps
The validation of automotive UIs traditionally relies on methodologies like specification-based testing [14] and Hardware-in-the-Loop (HiL) testing [15, 16]. Within HiL setups, visual assessment often employs Optical Character Recognition (OCR) and template matching [24, 32]. These conventional approaches are fragile, struggling with visual variations from updates or themes. Furthermore, they lack a deep semantic understanding of interface elements [30, 7]. Consequently, the maintenance overhead necessitates more intelligent solutions. This work bridges two critical gaps in prior research:
  1. 1.
    Automotive-Specific VLMs: Existing VLMs, including advanced models like the Molmo series with their spatial reasoning capabilities, are trained on generic UIs. This limits their generalization to automotive interfaces with custom iconography and layouts.
  2. 2.
    Evaluation-Centric Training: Unlike most LAMs that emphasize interaction (including Molmo’s pointing mechanism for interaction tasks), an evaluation-focused training approach is introduced in this research, enabling analysis and validation of automotive UIs.
Therefore, VLMs are leveraged to overcome the limitations of traditional methods, establishing a more robust and semantically aware framework for automotive UI validation. The development and evaluation of this framework constitute the core of this work. Specific recommendations for its integration into production-ready HiL testing environments are not provided.
@@ 3 Method
The methodology for adapting VLMs for automated automotive UI validation is outlined in this section. Domain adaptation challenges are addressed through an approach encompassing synthetic data generation and fine-tuning of baseline models, leading to the development of a specialized VLM for accurate visual grounding and evaluation of automotive UIs.
3.1 Domain Adaptation Challenges and Model Selection
Extensive research exists on LAMs and UI agents. However, these models primarily focus on interactions and implicitly derive evaluative functions by inferring necessary actions. For the targeted application of validating requirements via natural language test cases, explicit evaluation of expected outcomes is necessary. Existing smaller fine-tuned models (7B or less), such as TinyClick [27] and UGround [10], gather UI data from mobile, desktop, and web applications. This limits their applicability to automotive UIs. Domain-specific adaptation is required due to the distinct design patterns, custom iconography, and varied display hardware properties of modern automotive systems.
Specifically, the following challenges necessitated the proposed methodology:
@@ 3.1 Domain Adaptation Challenges and Model Selection
    Adapt VLMs for the domain of automotive UI: Modern automotive systems exhibit diverse UI designs influenced by brand, platform, driving mode, and display hardware properties. Icons pose a particular challenge, as little overlap exists with generic UIs from desktop, mobile, or web applications.
@@
    Assert evaluation capability: Explicit evaluation is not considered in prior work. To evaluate requirements or specifications, expected results must be prompted. These results require visual grounding and subsequent evaluation.
@@
    • Strict information security requirements are often held by Original Equipment Manufacturers. The leveraging of proprietary models such as OpenAI GPT-4o or Google Gemini Pro 2.0 presents challenges due to data privacy concerns and potential leakage of sensitive information to external servers.
    • Fine-tuning such models presents another challenge. While OpenAI and Google provide fine-tuning services for training and deployment, the details of the training process and the fine-tuned model itself remain closed source.
    • Challenging validation and prompt adaptation iterations to ensure intended functionality arise from the discontinuation of closed-source models.
    • The operation of models on local servers or even consumer-grade GPUs is highly desirable and necessary for a cost-effective solution.
To address these challenges, Evaluative Large Action Model (ELAM) is introduced, extending existing LAMs through the integration of explicit evaluation capabilities. The model is trained with synthetic data generated by an accompanying pipeline (Section 3.2). This pipeline is designed to address the unique characteristics of automotive UIs and the need for evaluative statements.
To establish a suitable baseline model for fine-tuning, four existing VLMs were evaluated: TinyClick (0.27B) [27], Molmo-7B-D-0924 [5], UGround-V1-7B (Qwen2-VL-based) [10], and InternVL2_5-8B [3]. TinyClick and UGround are specifically designed for UI interaction and point coordinate generation. However, their broader capabilities and suitability as a robust foundation for evaluation were limited. General-purpose VLMs, Molmo and InternVL2.5, have also been trained on UI data. The best performance (Table 2) among the considered models was achieved by Molmo-7B-D-0924 when evaluated on the proposed AutomotiveUI-Bench-4k dataset. Consequently, it was chosen as the foundation of ELAM-7B.
@@ 3.2 Synthetic Data Pipeline
Object-level captioning for automotive UIs presents significant challenges due to the need to incorporate functional, positional, and visual attributes to uniquely identify each element. Differentiating between Test Actions and Expected Results is required. Both necessitate object-level captions describing either the command initiating element actuation or a statement to be evaluated for the current UI. Furthermore, for Expected Results, a binary status (passed or failed) must be assigned, reflecting the validity of the expectation within the image. Empirical observation revealed that human annotators tend to under-represent failed expectations, likely due to the cognitive challenge of formulating incorrect expectations for a displayed UI image. It is also important to note that having human experts create data annotations is a time-consuming and expensive process. Therefore, a hybrid approach, leveraging human expertise for UI element selection via bounding box annotations combined with automated caption generation by large teacher models (e.g., GPT-4o, Gemini 2.0 Flash Thinking), offers a practical balance between annotation quality and time investment. Automated UI element extraction methods, such as OmniParser V1.5 [22] and a custom YOLOv8 [13] fine-tuned to automotive UI, were found to be insufficiently reliable, frequently resulting in misaligned bounding box detections that negatively impacted the correctness of generated captions.
The synthetic data pipeline (Figure 1) was used to train ELAM-7B. Bounding box regions from the dataset were prompted using the Set-of-Mark (SoM) technique [39]. Inspired by UGround [10], a bounding box indicated by an arrow was utilized. For Test Actions, prompts were designed as short yet comprehensive instructions that unambiguously map to a single UI element (Figure 2(a)). Prior to generating each Test Action, the model was instructed to provide a detailed reasoning of the process for identifying the target UI element, including its semantic, positional, and visual characteristics, as well as any relevant relationships to other elements, such as the current context or list headings (Figure 2). Utilized prompts can be found in Section A.3.
@@
Figure 2: Examples of synthetic data for training, illustrating (a) a test action “Tap the ’X’ button to cancel the displayed route guidance information on the map.” with prior reasoning “The requested element is an “X” icon located in the top right corner of the map display, semantically representing a close or cancel action, positioned next to the text “2 min” and “850 m,” which likely refers to route guidance information, with the parent element being the map view.” and (b) a failed expected result “The media control button in the bottom bar is expected to be a pause icon, indicating that media is currently playing. $\to\,$ Failed” with prior reasoning “The requested element is located in the bottom media control bar and is identified as a play symbol, which semantically represents the action to start or resume media playback. The evaluation highlights that there could be a misunderstanding regarding the icon’s function, as it may be incorrectly expected to represent a pause action instead. This confusion emphasizes the importance of clear iconography in user interface design to ensure users can easily understand the intended actions associated with each control.” in different infotainment systems.
To improve the generation of Expected Results, the process was modified to include a prior successful test action leading to the current menu or UI state (Figure 2(b)). Specifically, the model first generated a prior test action, followed by reasoning for the expectation and its “passed/failed” conclusion. Initially, this approach exhibited the same bias as human annotators and primarily generated passed validations. This issue was mitigated by explicitly prompting for evaluations that entail a failed result (Section A.3.3). After manually reviewing a small set of samples for GPT-4o, Gemini 2.0 Flash Thinking and Claude 3.5 Sonnet, Gemini (gemini-2.0-flash-thinking-exp-01-21) was selected to generate the training data. It was chosen for its better adherence to the required output structure, which was necessary for automatic parsing.
To mitigate potential overfitting on the specific linguistic style of the teacher model, the text for reasoning, test actions, and expectations were rephrased using a smaller model (gpt-4o-mini-2024-07-18). This rephrasing step is crucial for ensuring robustness and generalizability when processing inputs from real-world test case data. The teacher model, configured with a temperature of zero to minimize hallucinations, exhibited a tendency towards repetitive use of verbs and sentence structures, necessitating this diversification strategy. The contribution of rephrasing is explored in Table 4.